OpenStudy (anonymous):

If A and B are second quadrant angles such that cos A = -2/5 and sin B = 3/7, find the exact value of sin (A-B). Express your answer as one fraction; you do not need to rationalize.

OpenStudy (mertsj):

\[\sin(A-B)=sinAcosB-cosAsinB\]

OpenStudy (anonymous):

you still have to find \(\sin(B)\) and \(\cos(A)\) to finish this do you know how to do that?

OpenStudy (anonymous):

Yes, but plugging in the values into the equation would not give a final fraction, right? My teacher wants me to have in answer in fraction form.

OpenStudy (anonymous):

Okay never mind, I don't know what I'm doing now.

OpenStudy (mertsj):